OpenStudy (gudetamas):

Can someone check my answers? . Find the inverse of the given functions 1. g(x) = 15x - 10 answer: g-1(x)= 15x+ 2/3 2. j(x) = 3x+1/6 answer: j-1(x) = 2x - 1/6

OpenStudy (learner):

\(\color{blue}{\displaystyle f(x)=mx+b\quad \Longleftrightarrow \quad f^{-1}(x)=\frac{1}{m}x-\frac{\rm b}{m}}\)

OpenStudy (learner):

use that rule.

OpenStudy (gudetamas):

for one i ended up with g^-1(x) = x+10/15

OpenStudy (gudetamas):

and then i cross divided so it would be g-1(x)= 15x+ 2/3, im not sure how im wrong.